1. Suppose the utility function for goods q1 and q2 is given by

U(q1, q2) = q1q2 + q2

(a) Calculate the uncompensated (Marshallian) demand functions for q1 and q2

(b) Describe how the uncompensated demand curves for q1 and q2 are shifted by changes in

income (Y) or the price of the other good.

(c) Calculate the expenditure function for q1 and q2 such that

minimum expenditure = E(p1, p2, U)

(d) Use the expenditure function calculated in part (c) to compute the compensated demand

(Hicksian) functions for goods q1 and q2.

(e) Describe how the compensated demand curves for q1 and q2 are shifted by changes in

income (Y) or by changes in the price of the other good.
Answer #1

Suppose the utlity fuction for goods 21 and an Ř =) P2 Q2 = P, qi+P, -1) Pi (ait) in the is given byt U(91,92) = 9122792 of qHice, mas shallian or uncompusated demand fuctions for q, and as will be 9,- (Y-1) EPT 2P1 (It (*E+2A) 9+205-1) 2R and q2² 22Hunce as in come to incmans demand for good 2 for change in other goo price of otur goods? will incmase. da so 4 dan 2 will aE = 20/2/2/2 pil-pi NOW Partal i derivative of the enfudi tre fiction with respect to puces give is the U= q, a2 +22 =) U= 221/ a E a Pi = £ Pr? tice , hicksian demand fiction for good q1 e Now, compir sated demand for good al il Į pr? 20 P2/2_, » qeWhen income changes, Hice, as income changes, compusated demad When, price of other good chagus ! Now). On compusated demand